# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v4

@@ Appendix F Proof of Theorem 6
[…]
@@
[…]
@@
[…]
@@
[…]
@@
In the above inequality, we have used Taylor’s theorem in (a), Lemma 13 in (b) and $\varepsilon$ is a small number in (c).
@@ Appendix G Proof of Theorem 7
$\begin{split}&\|P_{n}^{(n)}-\delta_{0}^{n}\|_{w}^{*}\\ =&\|P_{n}^{(n)}(\bm{x})-\delta_{0}^{n}(\bm{x})\|_{w}^{*}\\ \leq&\|P_{n}^{(n)}(x_{1},\cdots,x_{n})-P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})\cdot\delta_{0}^{1}(x_{n})\|_{w}^{*}\\ +&\|P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})\cdot\delta_{0}^{1}(x_{n})-\delta_{0}^{n}(x_{1},\cdots,x_{n})\|_{w}^{*}.\\ \end{split}$ (176)
@@
Since $t_{n}$ is a scale in $\mathbb{R}$ and is independent with $\Psi(n)$, we have $-\frac{1}{2}\mu\Psi(n)\cdot t_{n}^{2}=O(\frac{\mu\cdot t_{n}^{2}}{n^{\tau}})\rightarrow 0$ as $n\rightarrow\infty$, which means (G) tends to 00 uniformly. Hence, we have $\|P_{n}^{(n)}(x_{1},\cdots,x_{n})-P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})\cdot\delta_{0}^{1}(x_{n})\|_{w}^{*}\rightarrow 0$.
@@
$\begin{split}&\|P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})-\delta_{0}^{n-1}(x_{1},\cdots,x_{n-1})\|_{w}^{*}\\ \leq&\|P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})-P_{n-2}^{(n)}(x_{1},\cdots,x_{n-2})\cdot\delta_{0}^{1}(x_{n-1})\|_{w}^{*}\\ +&\|P_{n-2}^{(n)}(x_{1},\cdots,x_{n-2})\cdot\delta_{0}^{1}(x_{n-1})-\delta_{0}^{n-1}(x_{1},\cdots,x_{n-1})\|_{w}^{*}.\\ \end{split}$ (180)
Note that two terms in the equation in (180) are $n-1$ dimension situation as (176). Using the technique repeatedly, finally we just need to prove $\|P_{1}^{(n)}(x_{1})-\delta_{0}^{1}(x_{1})\|_{w}^{*}\rightarrow 0$ as $n\rightarrow\infty$. The characteristic functions of $P_{1}^{(n)}(x_{1})$ and $\delta_{0}^{1}(x_{1})$ are $\exp\left(-\frac{1}{2}\mu\Psi(n)t_{1}^{2}\right)$ and $1$, respectively. Following the same arguments above, we have $\|P_{1}^{(n)}(x_{1})-\delta_{0}^{1}(x_{1})\|_{w}^{*}\rightarrow 0$.
@@ Appendix H Proof of Lemma 16
We just prove that $(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}$ approaches 00 and $1$ with different $\tau$ as $n\rightarrow\infty$, respectively. The logarithm of $(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}$ can be formulated as follows,
@@
  • When $\tau<\frac{1}{2}$, $01-2\cdot\tau>$ 0 and the above logarithm will approach $\infty$- ∞ as $n\rightarrow\infty$. Consequently, $(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}$ approaches 00 as $n\rightarrow\infty$ and the conclusion is obtained.
  • When $\tau>\frac{1}{2}$, it is obvious that $-\frac{1}{4}n^{1-2\tau}$ approaches 00 from the negative axis. Let $t=\left(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}}\right)^{\frac{n}{2}}$, from Taylor expansion, $e^{x}=1+x+o(x)$, we have
@@
    Thus, $H^{2}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ approaches 00 when $\tau>\frac{1}{2}$.

### v5

@@ Appendix F Proof of Theorem 8
$\begin{split}&\mathbb{V}(P_{1},P_{0})\\ =&\frac{1}{2}\int_{\mathbb{R}^{n}}|p_{1}(\bm{x})-p_{0}(\bm{x})|d\bm{x}\\ =&\frac{\varepsilon}{2}\int_{\mathbb{R}^{n}}p_{0}(\bm{x})|h_{\varepsilon}(\bm{x})|d\bm{x}\\ \leq&\frac{\varepsilon}{2}\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h_{\varepsilon}(\bm{x})|\\ =&\,\ \,\frac{1}{2}\varepsilon.\end{split}$ (184)
@@
$=$ $\int_{\mathbb{R}^{n}}p_{1}(\bm{x})\log\left(1+\varepsilon h_{\varepsilon}(\bm{x})\right)d\bm{x}$
$\overset{(a)}{=}$ $\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\left(1+\varepsilon h_{\varepsilon}(\bm{x})\right)\left[\epsilon h_{\varepsilon}(\bm{x})-\frac{1}{2}\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})\right]d\bm{x}$
@@
$=$ $\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\left[\varepsilon h_{\varepsilon}(\bm{x})+\frac{1}{2}\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})-\frac{1}{2}\varepsilon^{3}h_{\varepsilon}^{3}(\bm{x})\right]d\bm{x}$
@@
$\overset{(b)}{=}$ $\frac{1}{2}\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})d\bm{x}+o(\varepsilon^{2})$
@@
In the above inequality, we have used Taylor’s theorem in (a), Lemma 14 and the integrability assumption in (b) and $\varepsilon$ is a small number in (c).
@@ Appendix G Proof of Theorem 9
[…]
@@
Since $t_{n}$ is a scale in $\mathbb{R}$ and is independent with $\Psi(n)$, we have $\frac{1}{2}\mu\Psi(n)\cdot t_{n}^{2}=O(\frac{\mu\cdot t_{n}^{2}}{n^{\tau}})\rightarrow 0$ as $n\rightarrow\infty$, which means (G) tends to 0 uniformly. Hence, we have $\|P_{n}^{(n)}(x_{1},\cdots,x_{n})-P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})\cdot\delta_{0}^{1}(x_{n})\|_{w}^{*}\rightarrow 0$.
@@
[…]
Note that two terms in the equation in (190) are n1n-1 dimension situation as (186). Using the technique repeatedly, finally we just need to prove $\|P_{1}^{(n)}(x_{1})-\delta_{0}^{1}(x_{1})\|_{w}^{*}\rightarrow 0$ as $n\rightarrow\infty$. The characteristic functions of $P_{1}^{(n)}(x_{1})$ and $\delta_{0}^{1}(x_{1})$ are $\exp\left(-\frac{1}{2}\mu\Psi(n)t_{1}^{2}\right)$ and $1$, respectively. Following the same arguments above, we have $\|P_{1}^{(n)}(x_{1})-\delta_{0}^{1}(x_{1})\|_{w}^{*}\rightarrow 0$.
@@ Appendix H Proof of Lemma 17
We just prove that $(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}$ approaches 0 and $1$ with different $\tau$ as $n\rightarrow\infty$, respectively. The logarithm of $(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}$ can be formulated as follows,
@@
  • When $\tau<\frac{1}{2}$, $1-2\cdot\tau>0$ and the above logarithm will approach -$\infty$ as $n\rightarrow\infty$. Consequently, $(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}$ approaches 0 as $n\rightarrow\infty$ and the conclusion is obtained.
  • When $\tau>\frac{1}{2}$, it is obvious that $\frac{1}{4}n^{1-2\tau}$ approaches 0 from the negative axis. Let $t=\left(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}}\right)^{\frac{n}{2}}$, from Taylor expansion, $e^{x}=1+x+o(x)$, we have
@@
    Thus, $H^{2}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ approaches 0 when $\tau>\frac{1}{2}$.